# OpenVoice: Versatile Instant Voice Cloning

### 16 Dec 2023 v3

@@ 2.2 Model Structure
Although there are alternative ways [4, 9, 14] to extract $\mathbf{Z}(L_{I},S_{I})$, we empirically found that the proposed approach achieves the best audio quality. One can use HuBERT [4] to extract discrete or continuous acoustic units [14] to eliminate tone color information, but we found that such method also eliminates emotion and accent from the input speech. When the input is an unseen language, this type of method also has issues preserving the natural pronunciation of the phonemes. We also studied another approach [9] that carefully constructs information bottleneck to only preserve speech content, but we observed that this method is unable to completely eliminate the tone color.
@@ 3 Experiment
Accurate Tone Color Cloning. We build a test set of reference speakers selected from celebrities, game characters and anonymous individuals. The test set covers a wide voice distributions including both expressive unique voices and neutral samples in human voice distribution. With any of the 4 base speakers and any of the reference speaker, OpenVoice is able to accurately clone the reference tone color and generate speech in multiple languages and accents. We invite the readers to this website [^4] https://myshell-tts.vercel.app/accurate-tone-color-cloning for qualitative results.
Flexible Control on Voice Styles. A premise for the proposed framework to flexibly control the speech styles is that the tone color converter is able to only modify the tone color and preserves all other styles and voice properties. In order to confirm this, we use both our base speaker model and the Microsoft TTS with SSML to generate a speech corpus of 1K samples with diverse styles (emotion, accent, rhythm, pauses and intonation) as the base voices. After converting to the reference tone color, we observed that all styles are well-preserved. In rare cases, the emotion will be slightly neutralized, and one way that we found to solve this problem is to replace the tone color embedding vector of this particular sentence with the average vector of multiple sentences with different emotions from the same base speaker. This gives less emotion information to the flow layers so that they do not eliminate the emotion. Since the tone color converter is able to preserve all the styles from the base voice, controlling the voice styles becomes very straightforward by simply manipulating the base speaker TTS model. The qualitative results are publicly available on this website [^5] https://myshell-tts.vercel.app/flexible-voice-style-control.
Cross-Lingual Voice Clone with Ease. OpenVoice achieves near zero-shot cross-lingual voice cloning without using any massive-speaker data for an unseen language. It does require a base speaker of the language, which can be achieved with minimum difficulty with the off-the-shelf models and datasets. On our website [^6] https://myshell-tts.vercel.app/zero-shot-cross-lingual-voice-cloning, we provide an abundant of samples that demonstrates the cross-lingual voice clone capabilities of the proposed approach. The cross-lingual capabilities are two-fold:

### 21 Dec 2023 v4

@@ 2.2 Model Structure
Alternative Ways and Drawbacks. Although there are alternative ways [4, 9, 14] to extract $\mathbf{Z}(L_{I},S_{I})$, we empirically found that the proposed approach achieves the best audio quality. One can use HuBERT [4] to extract discrete or continuous acoustic units [14] to eliminate tone color information, but we found that such method also eliminates emotion and accent from the input speech. When the input is an unseen language, this type of method also has issues preserving the natural pronunciation of the phonemes. We also studied another approach [9] that carefully constructs information bottleneck to only preserve speech content, but we observed that this method is unable to completely eliminate the tone color.
Remark on Novelty. OpenVoice does not intend to invent the submodules in the model structure. Both the base speaker TTS model and the tone color converter borrow the model structure from existing work [5, 6]. The contribution of OpenVoice is the decoupled framework that seperates the voice style and language control from the tone color cloning. This is very simple, but very effective, especially when one wants to control styles, accents or generalize to new languages. If one wanted to have the same control on a coupled framework such as XTTS [3], it could require a tremendous amount of data and computing, and it is relatively hard to fluently speak every language. In OpenVoice, as long as the single-speaker TTS speaks fluently, the cloned voice will be fluent. Decoupling the generation of voice styles and language from the generation of tone color is the core philosophy of OpenVoice. We also provided our insights of using flow layers in tone color converter, and the importance of choosing a universal phoneme system in language generalization in our experiment section.
@@ 3 Experiment
Accurate Tone Color Cloning. We build a test set of reference speakers selected from celebrities, game characters and anonymous individuals. The test set covers a wide voice distributions including both expressive unique voices and neutral samples in human voice distribution. With any of the 4 base speakers and any of the reference speaker, OpenVoice is able to accurately clone the reference tone color and generate speech in multiple languages and accents. We invite the readers to this website [^4] https://research.myshell.ai/open-voice/accurate-tone-color-cloning for qualitative results.
Flexible Control on Voice Styles. A premise for the proposed framework to flexibly control the speech styles is that the tone color converter is able to only modify the tone color and preserves all other styles and voice properties. In order to confirm this, we use both our base speaker model and the Microsoft TTS with SSML to generate a speech corpus of 1K samples with diverse styles (emotion, accent, rhythm, pauses and intonation) as the base voices. After converting to the reference tone color, we observed that all styles are well-preserved. In rare cases, the emotion will be slightly neutralized, and one way that we found to solve this problem is to replace the tone color embedding vector of this particular sentence with the average vector of multiple sentences with different emotions from the same base speaker. This gives less emotion information to the flow layers so that they do not eliminate the emotion. Since the tone color converter is able to preserve all the styles from the base voice, controlling the voice styles becomes very straightforward by simply manipulating the base speaker TTS model. The qualitative results are publicly available on this website [^5] https://research.myshell.ai/open-voice/flexible-voice-style-control.
Cross-Lingual Voice Clone with Ease. OpenVoice achieves near zero-shot cross-lingual voice cloning without using any massive-speaker data for an unseen language. It does require a base speaker of the language, which can be achieved with minimum difficulty with the off-the-shelf models and datasets. On our website [^6] https://research.myshell.ai/open-voice/zero-shot-cross-lingual-voice-cloning, we provide an abundant of samples that demonstrates the cross-lingual voice clone capabilities of the proposed approach. The cross-lingual capabilities are two-fold: